Circular No. 7845

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002bl IN UGC 5499 M. Armstrong, Rolvenden, England, reports his discovery of an ap-

parent supernova (mag 17.0) on an unfiltered CCD images taken on Mar. 2.901, 5.879, and 7.059 in the course of the U.K. Nova/Supernova Patrol. The new object is located at α = 10^{h}12^{m}17^{s}.28, δ = +27^{o}51′52′′.4 (equinox 2000.0), which is ≈ 5′′ west and 9′′ north of UGC 5499. SN 2002bl is not present on Armstrong’s images from 2000 Dec. 22 (limiting mag 19.5) or 2002 Feb. 14 and Jan. 15, neither is it present on Palomar Sky Survey images taken on 1990 Jan. 1 (red, limiting mag 20.8) and 1993 Mar. 13 (blue, limiting mag 22.5).

A. V. Filippenko, University of California, Berkeley; D. C. Leonard, University of Massachusetts, Amherst; and E. C. Moran, Wesleyan University, write: “Inspection of a CCD spectrum (range 390–880 nm), obtained on Mar. 7 UT with the Keck-I telescope, reveals that the object is very similar to the peculiar type-Ic supernovae (‘hypernovae’) 1997ef, 1998bw, and 2002ap, ∼ 2 weeks after the explosion (Iwamoto et al. 2000, Ap.J. 534, 660, Fig. 1; see also IAUC 7825). The relative strength of the broad 535- and 645-nm emission features resembles that of SNe 1997ef and 2002ap more than that of SN 1998bw at any published phase (Patat et al. 2001, Ap.J. 555, 900, Fig. 2). On the other hand, the absorption trough nominally associated with the Ca II triplet is very weak, more closely resembling that of SN 1998bw than that of SNe 1997ef and 2002ap. The approximate expansion velocity derived from the broad, flat minimum of the O I 777.4-nm absorption trough is 17 000 km/s. The object is polarized at an average level of 0.4 percent; preliminary analysis suggests the presence of slight undulations that are correlated with features in the total flux spectrum.”

SUPERNOVA 2002bm IN MCG –01-32-19 T. Puckett, Mountain Town, GA; and A. Sehgal, Woodinville, WA,

report the discovery of an apparent supernova (mag 17.0) on an unfiltered CCD frame taken with the Puckett Observatory 0.35-m automated supernova patrol telescope on Mar. 6.24 UT. SN 2002bm is located at α = 12^{h}30^{m}16^{s}.90, δ = −8^{o}23′34′′.8 (equinox 2000.0), which is 11′′.4 east and 15′′.0 north of the center of MCG –01-32-19. The new object was confirmed on CCD frames taken on Mar. 7.33, but it is not present on images taken by Puckett on 2000 Nov. 30, Dec. 8, 2001 May 23, 2002 Jan. 16 (limiting mag ∼ 20.0) or on Palomar Sky Survey images taken on 1993 Jan. 24 (limiting mag ∼ 21.0) and 1984 Apr. 4 (limiting mag ∼ 20.0).

2002 March 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green